a collection of 75 coins consists of 25 cents and 50 cents. if the total value is 32.50 pesos, find the number of 25 cents and 50 cents coins in the collection.
Answer by mananth(16946)

x 25 cent coins
y 50 cent coins
Total tickets sold
1.00 x + 1.00 y = 75.00 .............1
Value of tickets
25.00 x + 50.00 y = 3,250.00 .............2
Eliminate y
multiply (1)by -50.00
Multiply (2) by 1.00
-50.00 x -50.00 y = -3,750.00
25.00 x + 50.00 y = 3,250.00
Add the two equations
-25.00 x = -500.00
/ -25.00
x = 20.00
plug value of x in (1)
1.00 x + 1.00 y = 75.00
20.00 + y = 75.00
y = 75.00 -20
y = 55.00
y = 55.00
20 25 cent coins
55 50 cent coins
m.ananth@hotmail.ca